The wind chill factor represents the equivalent air temperature at a standard wind speed that would produce the same heat loss as the given temperature and wind speed. One formula for computing the equivalent temperature is


where v represents the wind speed (in meters per second) and t represents the air temperature . Compute the wind chill for an air temperature of 15°C and a wind speed of 12 meters per second. (Round the answer to one decimal place.) Show your work.

W(t) = {33-(10.45+10sqrtv-v)(33-t)/2204
33-1.5958(33-t)
if 0≤v<1.79
if 1.79≤v<20
if v≥20



Answer by Fombitz(32388)

You need to rewrite your equation, it's all run together.
Make separate lines for each segment.
As an example,
f(x)=x+1 for x<0
f(x)=x^2+2 for 0<=x<=2
f(x)=-x for x>2
.
.
.
Please repost.